OpenStudy (anonymous):

At the park, Joanne and Maria bought food and lemonade for themselves and their friend, Brian. Joanne bought 3 slices of pizza and 1 lemonade for $4.50. Maria bought 2 slices of pizza and 2 lemonades (one for her and one for Brian) for $4.00. How much does Brian owe Maria for his lemonade?

OpenStudy (anonymous):

let p = the cost of one pizza slice. let l = the cost of one lemonade So we have: 3p + 1l = 4.50 2p + 2l = 4.00 multiply the top equation by two multiple the bottom equation by -1 6p + 2l = 9.00 -2p - 2l = -4.00 4p = 5.00 p= 1.25 Substitute 1.25 for p in the top equation. 3*1.25 +1l = 4.50 1l = 4.50 -3.75 l= .75 So one lemonade costs 75 cents.